ninajaen:

Multiply (3)(−4)(2).

jhonyy9:

ok step by step (3)×(-4) = ?

jhonyy9:

but you need to know 3 is +3 and when multiplie plus by minus what result ?

ninajaen:

is it -12

jhonyy9:

perfect so (-12)×(+2)= ?

Lura:

(U might know already, might not even be related) But, just in case, Positive and negatives or vice versa, are still negative unless both are positive or negative.

jhonyy9:

@ninajaen courage !!! 12×2 = ?

ninajaen:

24

jhonyy9:

yes - but there is -12 ×2 so in this case will be = ?

ninajaen:

-24

Lura:

good job ^^

jhonyy9:

perfect

ninajaen:

:) thank you

jhonyy9:

np any time

Lura:

yw my guy :>